OpenStudy (anonymous):

serious help needed never solved a word problem using algebra before. Your little brother wants to set up a juice stand this weekend. It is going to cost £7 to get everything set up and then it will cost 15 pence per cup to make. Your brother wants to sell the juice for 50 pence per cup. How many cups will have to be sold in order to break even.

OpenStudy (anonymous):

I don't know how the euro works (in terms of currency). But I assume a Euro is similar to a dollar, and pences are like "cents"

OpenStudy (anonymous):

Ahh I'll just use dollars and cents.

OpenStudy (anonymous):

So $7 to set up , $0.15 per cup and we are selling it for $0.50 per cup

OpenStudy (anonymous):

Total cost = Fixed Cost +Variable Cost. y= 7 + 0.15x Where y= Total Cost and x=number of cups sold.

OpenStudy (anonymous):

Total revenue= 0.5x ( 50cents per cup, and you sell "x" cups , so 0.5*x = 0.5x)

OpenStudy (anonymous):

Equate them and solve 7+0.15x = 0.5x x= 7/0.35 = 20 So you must sell 20cups to break even.

OpenStudy (anonymous):

cheers